c=5/9(f-32)for f (fahrenheit to celsius)
Answer by stanbon(75887)

c=5/9(f-32)for f
-----------
C = (5/9)(f-32)
---
f-32 = (9/5)C
---
F = (9/5)C + 32
===================
Cheers,
Stan H.